A solid cylinder of mass 20 kg, has length 1 metre and radius 0.5m. then its momentum of inertia in kg m2 about its geometrical axis is
(a) 2.5 (b) 5
(c) 1.5 (d) 3

Grade:12

1 Answers

Vikas TU
3 years ago
Mass of cylinder (M) = 20 Kg
Radius of its ( R) = 0.25 m
Angular speed (w ) = 100 rad/s
We know, moment of inertia of solid cylinder ( I) = (1/2)MR²
= (1/2) × 20× (0.25)²
= 0.625 kg.m²
Kinetic energy associated with the rotation of the cylinder is given by
K.E = (1/2)Iw²
= (1/2) × 0.625 × (100)²
= 3125 joule

angular momentum = Iw = 0.625* 100 = 62.5 js